The car A is moving under the conditions shown on a circular road. Evaluate its velocity and acceleration magnitudes. Ans: a 76.84 m/s2 r 300 ft 6 0.4 rad/s e 0.2 rad/s2
Answer #1

Velocity = r * w

Given

r = 300 ft = 91.44 m

w = angular velocity = 0.4 rad/ s2

Therefore, linear velocity = v = r* w = 91.44 * 0.4 = 36.57 m/s

centripetal acceleration = r* w​​​​​2 = 91.44* 0.42 = 14.63 m/s2

Tangential acceleration = alpha*r = 0.2* 91.44= 18.28 m/ s2